OpenStudy (anonymous):

Find the slope-intercept form of this equation... 3x-2y=-4

OpenStudy (anonymous):

\[Slope~Intercept~Form\] \[y = mx + b\]

OpenStudy (anonymous):

Where: \[m~is~the~slope\] \[b~is~the~y-intercept\]

OpenStudy (anonymous):

3x-2y=-4 Solve for y. 3x-2y=-4 *subtract 3x from both sides* 3x-2y=-4 -3x -3x ---- ----- -2y = -3x - 4 *Divide by -2 on both sides* -2y = -3x - 4 ---- --- --- -2 -2 -2 *what do you get?*

OpenStudy (jhannybean):

Your equation is in standard form \[\large Ax+By = C\] All you need to so is change it to slpe intercept form \[\large y=mx+b\] where b= y-intercept and m = slope.

OpenStudy (anonymous):

y= 1.5x +2

OpenStudy (anonymous):

\[y = \frac{ 3x }{ 2 } + 2\]

OpenStudy (anonymous):

seems right.

OpenStudy (jhannybean):

\[\large y=\color{blue}{1.5}x+\color{green}3\]\[\large \color{blue}{m=\text{slope}} \ \, \ \color{green}{b= \text{y-intercept}}\]

OpenStudy (anonymous):

+3?

OpenStudy (jhannybean):

ohh +2***

OpenStudy (jhannybean):

-_- typo.

OpenStudy (anonymous):

lol Jhanny

OpenStudy (anonymous):

-_-

OpenStudy (jhannybean):

Sorry!!

OpenStudy (anonymous):

ok back to business guys ^_^

OpenStudy (anonymous):

lol its okay :)

OpenStudy (anonymous):

so what next?

OpenStudy (anonymous):

umm... idk.. ur da one with the open question lol

OpenStudy (anonymous):

i mean in solving...

OpenStudy (anonymous):

what do you mean? You just said: Find the slope-intercept form of this equation... 3x-2y=-4 and we did...

OpenStudy (anonymous):

so what is the slope?

OpenStudy (anonymous):

Find the slope-intercept form of this equation... 3x-2y=-4 Point Slope Form of 3x-2y=-4 is y = 3/2x + 2 Point slope formula: y = mx + b So: Slope = m = 3/2 y-intercept = b = 2

OpenStudy (jhannybean):

I stated what the slope was... i just made a typo in typing a 3 instead of a 2 for the y-intercept.... <_<

OpenStudy (anonymous):

lol yep... ^ Jhanny did :)

OpenStudy (anonymous):

so how would i draw this?

OpenStudy (anonymous):

please help @some_someone

OpenStudy (anonymous):

well its a line.

OpenStudy (anonymous):

do you know how to draw lines?

OpenStudy (anonymous):

a line from where to where?

OpenStudy (jhannybean):

Start at the y-intercept. Y-intercept = 2

OpenStudy (jhannybean):

Now we take our slope which was 3/2, and from this point (0,2), we move up 3, and over 2.

OpenStudy (jhannybean):